OpenStudy (anonymous):

Find the distance between the pair of parallel lines, y = -2x+1 & y = -2x+16. Choose one answer. a. square root of 45, or appx 6.7 b. 15 c. square root of 54, or appx 7.3 d. square root of 50, or appx 7.1

OpenStudy (anonymous):

Step #1: determine an equation of a perpendicular line to these 2 lines. Step #2: solve 2 linear systems. The first is between the new line and the first of the parallel lines. The second is between the new line and the second of the parallel lines. Step #3: use distance formula on the two points determined from step #2.

OpenStudy (anonymous):

That didn't help much. Idk how to do this

OpenStudy (anonymous):

To start off with step #1, you'll know the slope of the perpendicular since it is the negative reciprocal of the given lines.

OpenStudy (anonymous):

lol can u just give me the answer?

OpenStudy (anonymous):

We have to show you HOW to get the answer at this website and guide you through the process. That process is done with your efforts. This website isn't an "ask questions, get answers with no explanations and learn nothing" website. You won't learn anything and you'll just get stuck always asking questions for every problem you have. Plus, students and tutors get in trouble for just giving answers. Just giving answers will not help anyway.

OpenStudy (anonymous):

Show how u get the answer thats how I learn

OpenStudy (anonymous):

that'll be fine

OpenStudy (anonymous):

Idk how to do this

OpenStudy (anonymous):

np. So, for the perpendicular line slope, it will be the negative reciprocal of -2 which is +1/2 or 1/2 because (-2)(1/2) = -1 and we got -2 from the original equations that are in the form y= mx + b. For the equation of the perpendicular line, we can use the origin, which is point (0, 0), for convenience. The point-slope form of the equation is:\[y - y _{1} = m(x - x _{1})\]where (x1, y1) is (0, 0) and "m" = slope =1/2, so we get y - 0 = (1/2)(x - 0) which is y = x/2 end of step #1

OpenStudy (anonymous):

Ok I understaNd that part

OpenStudy (anonymous):

Step #2: part a) solve the linear system: y = x/2 and y = -2x + 1. substituting: x/2 = -2x + 1, so x = -4x + 2 x = 2/5 so y = 1/5 part b) solve the linear system: y = x/2 and y = -2x + 16. susbstituting: x/2 = -2x + 16, so x = -4x + 32 x = 32/5 so y = 16/5 end of step #2

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

step #3: distance formula: d = sqrt[(x1 - x2)^2 + (y1 - y2)^2] d = sqrt[(32/5 - 2/5)^2 + (16/5 - 1/5)^2] at this point, it's just simplification. Still need help?

OpenStudy (anonymous):

Yeah. Im not good at math at all

OpenStudy (anonymous):

np. d = sqrt[(30/5)^2 + (15/5)^2] = sqrt[(6)^2 + (3)^2] =sqrt(36 + 9) = sqrt(45) and that's it! Yes, it's a lot of steps, but they are all logical and if you break it down this way, you'll be ok. The hardest part is conceptualizing the perpendicular line as the shortest distance between the two lines.

OpenStudy (anonymous):

Thank you. I actually understand it. Ur better than my geometry teacher

OpenStudy (anonymous):

math can be tough or you can be tough on math. I try to explain because students eventuright away like that better. Gives them independence.

OpenStudy (anonymous):

And you're quite welcome.